# On the Curses of Future and History in Future-dependent Value Functions for Off-policy Evaluation

### v1

@@ B.1 Proof of Lemma 1
B.2 Proof of Theorem 2
The proof follows similar idea from Uehara et al. (2022). Let $\mathcal{G}_{h}V=\mu(o_{h},a_{h})(r_{h}+V(f_{h+1}))-V(f_{h})$, we have
@@ B.2 Proof of Theorem 2
The second inequality is from Equation 16 and the last inequality is from Section B.2. We then have
@@
Combining Section B.2 and Equation 17, we have
$\left|\sum_{h=1}^{H}\{\mathbb{E}_{\mathcal{D}}-\mathbb{E}_{\pi_{b}}\}[(\xi(\tau_{h})-\mathcal{G}_{h}V)^{2}-((\mathcal{B}^{\mathcal{H}}V)(\tau_{h})-\mathcal{G}_{h}V)^{2}]\right|\leq\varepsilon_{\textrm{stat}}.$ (18)
@@
$\leq\sum_{h=1}^{H}\left|(\mathbb{E}_{\mathcal{D}}-\mathbb{E}_{\pi_{b}})[(\mathcal{G}_{h}\widehat{V})^{2}-(\mathcal{G}_{h}\widehat{V}-(\mathcal{B}^{\mathcal{H}}\widehat{V})(\tau_{h}))^{2}]\right|$
@@
Appendix C Proofs for Section 4
C.1 Example 1
@@ C.4 Proof of Proposition 4
C.5 Example 3
We give the calculation for Example 3. When $\pi_{e}=\pi_{b}$, $V_{{\mathcal{S}},h}^{\pi_{e}}=M_{\mathcal{F},h}R_{h}$, so
$~{}\|V_{{\mathcal{S}},h}^{\pi_{e}}\|_{\Sigma_{\mathcal{F},h}^{-1}}^{2}=R_{h}^{\top}M_{\mathcal{F},h}^{\top}\Sigma_{\mathcal{F},h}^{-1}M_{\mathcal{F},h}R_{h}$
$=$ $~{}R_{h}^{\top}Z_{h}^{1/2}Z_{h}^{-1/2}M_{\mathcal{F},h}^{\top}(M_{\mathcal{F},h}Z_{h}^{-1}M_{\mathcal{F},h}^{\top})M_{\mathcal{F},h}Z_{h}^{-1/2}Z_{h}^{1/2}R_{h}$
$\leq$ $~{}R_{h}^{\top}Z_{h}^{1/2}Z_{h}^{1/2}R_{h},$
@@ C.5 Example 3
$R_{h}^{\top}Z_{h}R_{h}=S\cdot\mathbb{E}_{Z_{h}/S}[R_{h}^{2}]\leq SH^{2}.$
Appendix D Proofs for Section 5
D.1 Proof of Lemma 5
@@
The last step follows from 10. ∎
D.2 Example 5
@@
D.3 Proof of Proposition 6
@@ D.6 Proof of Lemma 8
$=R(f_{h})\left\|\frac{\mathbf{u}(f_{h})}{Z(f_{h})}\right\|_{1}\left\|(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}\right\|_{\infty}$
$\leq R(f_{h})C_{\mathcal{F},\infty}$
@@
The second inequality is from 12 and $\frac{\mathbf{u}(f_{h})}{Z(f_{h})}$ is a stochastic vector. The last step is from the boundedness of the reward. ∎
D.7 Example 8
@@
$\Sigma_{\mathcal{F},h}^{R}\mathbf{1}=M_{\mathcal{F},h}(Z_{h}^{R})^{-1}M_{\mathcal{F},h}^{\top}\mathbf{1}=M_{\mathcal{F},h}(Z_{h}^{R})^{-1}\textrm{diag}(Z_{h})^{\top}=M_{\mathcal{F},h}R_{h}=V_{{\mathcal{S}}}^{\pi_{e}}.$
D.8 Example 9
The calculation is similar to that of Example 2 in Appendix C.3, except that the numerator has an extra $R(f_{h})$. Therefore, when calculating the $i$-th diagonal entry of $\Sigma_{\mathcal{F},h}^{R}$, the final sum is calculating the expectation of $R(f_{h})$ conditioned on $s_{h}=i$ under policy $\pi_{b}$, which is the definition of $V_{{\mathcal{S}}}^{\pi_{b}}(i)$.
D.9 Proof of Theorem 9
@@
D.10 Proof of Theorem 10
Finally, we prove Theorem 10. The proof uses the similar idea from Xie & Jiang (2020). For any $V\in\mathcal{V}$ and $w\in\mathcal{W}$, we define the population loss estimator $\mathcal{L}_{d^{\pi_{b}}}$ and the empirical loss estimator $\mathcal{L}_{\mathcal{D}}$ as follows
@@ D.10 Proof of Theorem 10
According to Lemma 8, 11, $C_{\mathcal{V}}\leq c\|V_{\mathcal{F}}\|_{\infty}$ and $C_{\mathcal{W}}\leq c\|w^{\star}\|_{\infty}$, we further have

### v2

@@ C.1 Proof of Lemma 1
C.2 Proof of Theorem 2
The proof follows similar idea from Uehara et al. [2022a]. Let $\mathcal{G}_{h}V=\mu(o_{h},a_{h})(r_{h}+V(f_{h+1}))-V(f_{h})$, we have
@@ C.2 Proof of Theorem 2
The second inequality is from Equation 12 and the last inequality is from Section C.2. We then have
@@
Combining Section C.2 and Equation 13, we have
$\left|\sum_{h=1}^{H}\{\mathbb{E}_{\mathcal{D}}-\mathbb{E}_{\pi_{b}}\}[(\widehat{\xi}_{V}(\tau_{h})-\mathcal{G}_{h}V)^{2}-((\mathcal{B}^{\mathcal{H}}V)(\tau_{h})-\mathcal{G}_{h}V)^{2}]\right|\leq\varepsilon_{\textrm{stat}}.$ (14)
@@
$\leq\left|\sum_{h=1}^{H}(\mathbb{E}_{\mathcal{D}}-\mathbb{E}_{\pi_{b}})[(\mathcal{G}_{h}\widehat{V})^{2}-(\mathcal{G}_{h}\widehat{V}-(\mathcal{B}^{\mathcal{H}}\widehat{V})(\tau_{h}))^{2}]\right|$
@@
Appendix D Proofs for Section 4
D.1 Example 1
@@ D.4 Proof of Proposition 4
D.5 Example 3
We give the calculation for Example 3. When $\pi_{e}=\pi_{b}$, $V_{{\mathcal{S}},h}^{\pi_{e}}=M_{\mathcal{F},h}R_{h}^{+}$, so
$~{}\|V_{{\mathcal{S}},h}^{\pi_{e}}\|_{\Sigma_{\mathcal{F},h}^{-1}}^{2}=(R_{h}^{+})^{\top}M_{\mathcal{F},h}^{\top}\Sigma_{\mathcal{F},h}^{-1}M_{\mathcal{F},h}R_{h}^{+}$
$=$ $~{}(R_{h}^{+})^{\top}Z_{h}^{1/2}Z_{h}^{-1/2}M_{\mathcal{F},h}^{\top}(M_{\mathcal{F},h}Z_{h}^{-1}M_{\mathcal{F},h}^{\top})M_{\mathcal{F},h}Z_{h}^{-1/2}Z_{h}^{1/2}R_{h}^{+}$
$\leq$ $~{}(R_{h}^{+})^{\top}Z_{h}^{1/2}Z_{h}^{1/2}R_{h}^{+},$
@@ D.5 Example 3
$(R_{h}^{+})^{\top}Z_{h}R_{h}^{+}=S\cdot\mathbb{E}_{Z_{h}/S}[(R_{h}^{+})^{2}]\leq SH^{2}.$
Appendix E Proofs for Section 5
E.1 Proof of Lemma 6
@@
The last step follows from 11. ∎
E.2 Example 5
@@
E.3 Comparison between belief coverage and latent state coverage
Here we show that belief coverage is stronger than latent state coverage. More concretely, consider a standard measure of latent state coverage, the 2nd moment of state density ratio (c.f. discussion below Example 7):
$\mathbb{E}_{\pi_{b}}[(d^{\pi_{e}}(s_{h})/d^{\pi_{b}}(s_{h}))^{2}]=(\mathbf{b}_{h}^{\pi_{e}})^{\top}\textrm{diag}(\mathbb{E}_{\pi_{b}}[\mathbf{b}(\tau_{h})])^{-1}\mathbf{b}_{h}^{\pi_{e}}.$
In comparison, our belief coverage parameter from 11 is
$(\mathbf{b}_{h}^{\pi_{e}})^{\top}\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}=(\mathbf{b}_{h}^{\pi_{e}})^{\top}\mathbb{E}_{\pi_{b}}[\mathbf{b}(\tau_{h})\mathbf{b}(\tau_{h})^{\top}]^{-1}\mathbf{b}_{h}^{\pi_{e}}.$
To show the former is smaller than the latter, it suffices to show that
$\textrm{diag}(\mathbb{E}_{\pi_{b}}[\mathbf{b}(\tau_{h})])^{-1}\preceq(\mathbb{E}_{\pi_{b}}[\mathbf{b}(\tau_{h})\mathbf{b}(\tau_{h})^{\top}])^{-1},$
which is implied by $\textrm{diag}(\mathbb{E}_{\pi_{b}}[\mathbf{b}(\tau_{h})])\succeq\mathbb{E}_{\pi_{b}}[\mathbf{b}(\tau_{h})\mathbf{b}(\tau_{h})^{\top}]$. It therefore suffices to show that $\textrm{diag}(\mathbf{b}(\tau_{h}))\succeq\mathbf{b}(\tau_{h})\mathbf{b}(\tau_{h})^{\top}$ holds in a pointwise manner for all $\tau_{h}$. To show this, we temporarily let $\mathbf{b}=\mathbf{b}(\tau_{h})$ in the calculation: consider an arbitrary vector $v\in\mathbb{R}^{S}$, then
$v^{\top}(\textrm{diag}(\mathbf{b})-\mathbf{b}\mathbf{b}^{\top})v=\mathbb{E}_{s\sim\mathbf{b}}[v(s)^{2}]-(\mathbb{E}_{s\sim\mathbf{b}}[v(s)])^{2}\geq 0.$
The last step is due to Jensen’s inequality.
E.4 Comparison to $\textrm{IV}(\mathcal{V})$
We now compare to the $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}$ terms in Theorem 2. Belief coverage is generally stronger than latent state coverage which corresponds to the $\textrm{Dr}_{\mathcal{V}}$ term (see Section E.3). That said, perhaps surprisingly, the remaining $\textrm{IV}(\mathcal{V})$ term must scale with $\frac{1}{\sigma_{\min}(\Sigma_{\mathcal{H},h})}$ under moderate assumptions.
Proposition 10.
The proposition states that as long as $\min_{s_{h}}d^{\pi_{b}}(s_{h})$ is bounded away from 00 (which is benign and helps bound the $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ term) and $\mathcal{V}$ is sufficiently rich that $\mathcal{B}^{{\mathcal{S}}}_{h}\mathcal{V}$ includes a certain vector in $\mathbb{R}^{S}$, then boundedness of $\textrm{IV}(\mathcal{V})$ requires bounded $1/\sigma_{\min}(\Sigma_{\mathcal{H},h})$, which is stronger than our belief coverage assumption. This renders the split between $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ superficial and perhaps unnecessary.
Proof of Proposition 10.
@@ E.7 Proof of Lemma 5
$=R^{+}(f_{h})\left\|\frac{\mathbf{u}(f_{h})}{Z(f_{h})}\right\|_{1}\left\|(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}\right\|_{\infty}$
$\leq R^{+}(f_{h})C_{\mathcal{F},\infty}$
@@
The second inequality is from 9 and $\frac{\mathbf{u}(f_{h})}{Z(f_{h})}$ is a stochastic vector. The last step is from the boundedness of the reward. ∎
E.8 Example 4
@@
$\Sigma_{\mathcal{F},h}^{R}\mathbf{1}=M_{\mathcal{F},h}(Z_{h}^{R})^{-1}M_{\mathcal{F},h}^{\top}\mathbf{1}=M_{\mathcal{F},h}(Z_{h}^{R})^{-1}\textrm{diag}(Z_{h})^{\top}=M_{\mathcal{F},h}R_{h}^{+}=V_{{\mathcal{S}}}^{\pi_{e}}.$
E.9 Example 9
The calculation is similar to that of Example 2 in Appendix D.3, except that the numerator has an extra $R^{+}(f_{h})$. Therefore, when calculating the $i$-th diagonal entry of $\Sigma_{\mathcal{F},h}^{R}$, the final sum is calculating the expectation of $R^{+}(f_{h})$ conditioned on $s_{h}=i$ under policy $\pi_{b}$, which is the definition of $V_{{\mathcal{S}}}^{\pi_{b}}(i)$.
E.10 Proof of Theorem 7
@@
E.11 Proof of Theorem 9
Finally, we prove Theorem 9. The proof uses the similar idea from Xie and Jiang [2020]. For any $V\in\mathcal{V}$ and $w\in\mathcal{W}$, we define the population loss estimator $\mathcal{L}_{d^{\pi_{b}}}$ and the empirical loss estimator $\mathcal{L}_{\mathcal{D}}$ as follows
@@ E.11 Proof of Theorem 9
According to Lemma 5, 12, $C_{\mathcal{V}}\leq c\|V_{\mathcal{F}}\|_{\infty}$ and $C_{\mathcal{W}}\leq c\|w^{\star}\|_{\infty}$, we further have